Karlottalotta:

Point E is located at (2, −3) and point F is located at (−2, −1). Find the y value for the point that is 3 over 4 the distance from point E to point F. (5 points) −2.5 −4.5 −3.5 −1.5

MxxnLight:

Any thoughts?

XxAngelGirlxX:

E (2, -3) F (-2, -1) the y distance is from -3 to -1 which is a difference of 2 2(3/4) = 1 1/2 or 1.5 -3 + 3/2 ?

Sweetie145:

3/4 the distance from point e to point f means that the point divides the line ef in the ratio 3:1 x = (3(-2) + 1(2))/(3 + 1) = (-6 + 2)/4 = -4/4 = -1 y = (3(-1) + 1(-3))/(3 + 1) = (-3 - 3)/4 = -6/4 = -1.5 required point = (-1, -1.5)